OpenStudy (anonymous):

HELPPPPPPPPPPPPP PPPPPPPPLLLLLLLLLLEEEASEEEEE Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3

OpenStudy (anonymous):

y - 6 = -3x - 3 y - 6 = -3(x + 1) y = -3x + 3 3x + y = 3